Chapman Company obtains 100 percent of Abernethy Company’s stock on January 1, 2017. As of that date, Abernethy has the following trial balance:

Debit Credit
Accounts payable $ 52,400
Accounts receivable $ 48,600
Additional paid-in capital 50,000
Buildings (net) (4-year remaining life) 179,000
Cash and short-term investments 61,250
Common stock 250,000
Equipment (net) (5-year remaining life) 260,000
Inventory 121,500
Land 105,000
Long-term liabilities (mature 12/31/20) 174,500
Retained earnings, 1/1/17 264,650
Supplies 16,200
Totals $ 791,550 $ 791,550

During 2017, Abernethy reported net income of $86,000 while declaring and paying dividends of $11,000. During 2018, Abernethy reported net income of $124,500 while declaring and paying dividends of $47,000.

Assume that Chapman Company acquired Abernethy’s common stock for $675,160 in cash. Assume that the equipment and long-term liabilities had fair values of $284,350 and $142,140, respectively, on the acquisition date. Chapman uses the initial value method to account for its investment.

Prepare consolidation worksheet entries for December 31, 2017, and December 31, 2018.
Answer #1
  • Calculations and Entries for 31.12.2017

1.Total Stockholders Equity = Additional Paid in capital + Common Stock + Retained Earnings

= 50,000 + 250,000 + 264,650 = 564,650

2.Difference of Acquisition Price and Total Stockholders Equity = 675,160 - 564650 = 110,510

3.Equipment Fair Value = 284,350 Book Value = 260,000

Difference = 284,350 - 260,000 = 24,350

4. Long term Liability Fair Value = 142,140 Book Value = 174,500

Difference = 174500 - 142140 = 32,360

5. Excess of Fair Value over Book Value = 24,350 + 32,360 = 56,710

6. Goodwill = 110,510 - 56710 = 53,800

7. Amortization of Equipment = 24,350/5 = 4,870

8. Amortization of Long Term Liability = 32360/4 = 8,090

9. Total Amortization = 4,870 + 8,090 = 12,960

10. Equity Income = Net Income - Amortization = 86,000 - 12,960 = 73,040

  • Journal Entries
Particulars Amount ($) Amount ($)

Common Stock.....Dr.

250,000
Additional Paid In Capital.....Dr. 50,000
Retained Earnings.....Dr. 264,650
To Investment in Company A 564,650
Equipment....Dr. 24,350
Long Term Liability....Dr. 32,360
Goodwill....Dr. 53,800
To Investment in Company A 110,510
Equity Earnings In Company A....Dr. 73,040
To Investment in Company A 73,040
Dividend Income....Dr. 11,000
To Dividend Paid 11,000
Depreciation.....Dr. 4,870
Interest.....Dr. 8,090
To Equipment 4,870
To Long Term Liability 8,090
  • Calculations and Entries for 31.12.2018

1. Retained Earnings for Company C - Dividends - Amortization = 86,000 - 11,000 - 12,960 = 62,040

​​​​​​​ 2. Retained Earnings for Company C at the beginning of 2018

= Retained Earnings on 1.1.17 + Net Income of 2017 - Dividends

= 264,650 + 86,000 - 11,000 = 339,650

3. Total Investment in Company A = 564,650 + 86,000 - 11,000 = 639,650

4. Balance of Total Fair Value over Book Value - Amortization = 110,510 - 12,960 = 97,550

5. Balance of Fair Value over Book Value for Equipment - Amortization = 24,350 - 4,870 = 19,480

6. Balance of Fair Value over Book Value for Long Term Liability - Amortization = 32,360 - 8,090 = 24,270

  • Journal Entries
Particulars Amount ($) Amount ($)
Investment in Company A.....Dr. 62,040
To Retained Earnings on 1.1.18 62,040
Common Stock.....Dr. 250,000
Additional Paid In Capital.....Dr. 50,000
Retained Earnings......Dr. 339,650
To Investment in Company A 639,650
Equipment.....Dr. 19,480
Long Term Liability.....Dr. 24,270
Goodwill.....Dr. 53,800
To Investment in Company A 97,550
Dividend Income....Dr. 47,000
To Dividend Paid 47,000
Depreciation.....Dr. 4,870
Interest.....Dr. 8,090
To Equipment 4,870
To Long Term Liability 8,090